Grade 11Algebra

ln2x^ln2=ln3y^ln3
3^lnx=2^lny
then find the values of x and y

sri priya yadav
7 Years agoGrade 11

2 Answers

kartik
7 Years ago
(2x)^ln2 = (3y)^ln3
3^lnx= 2^lny
taking log on both sides
ln2 ln2x = ln3 ln3y                                                                                                      lnx ln3= lny ln2 
using the 2 equations eleminate one variable and solve for other
Suraj
5 Years ago
(2x) 
ln2
 =(3y) 
ln3
 
Taking ln both side,
⇒(ln2)ln(2x)=(ln3)ln(3y)
⇒(ln2)(ln2+lnx)=(ln3)(ln3+lny)
⇒ 
(lny)
(ln2)
 (ln2+lnx)=(ln3)(1+ 
lny
ln3
 )
⇒(ln2)( 
lny
ln2
 + 
lny
lnx
 )=ln3(1+ 
lny
ln3
 )(equation 1)
And 3 
lnx
 =2 
lny
 
Taking ln both side,
⇒(lnx)(ln3)=(lny)(ln2)
⇒ 
lny
lnx
 = 
ln3
ln2
 (equation 2)
From 1 and 2
⇒(ln2)( 
lny
ln2
 + 
lny
lnx
 )=ln3(1+ 
lny
ln3
 )
⇒ 
lny
(ln2) 
2
 −(ln3) 
2
 
 = 
ln3
(ln3) 
2
 (ln2) 
2
 
 
⇒lny=−ln3
⇒y= 
3
1
 
And from 1 
⇒lnx=−ln2
⇒x=+ 
2
1